OpenStudy (anonymous):

cos(x-y) when cosx=3/5 and cos y=4/5 Last problem before I don't understand. Exam in four hours =/ Please help :D

OpenStudy (anonymous):

You're supposed to find the exact value. Cos(x-y)=cosx cosy + sinx siny

OpenStudy (anonymous):

cosx=3/5 >>>cos^2x=(9/25) sin^2x+cos^2x=1>>>> sin^2x=1-cos^2x>>> sinx = 4/5 >>>>>> on the sameway you can found sin y go on , cosx=cosx*cosy+sinx siny >>>

OpenStudy (anonymous):

get it